# AI Scaling: From Up to Down and Out

### v1

@@ 1 Introduction
The field of artificial intelligence (AI) has witnessed extraordinary advancements over the past decade, largely driven by the relentless pursuit of Scaling Up. Early breakthroughs were characterized by models with millions of parameters, such as AlexNet (Krizhevsky et al., 2012), word2vec (Church, 2017) and BERT (Devlin, 2018), which paved the way for deep learning’s success. This progression quickly escalated to models with billions of parameters, exemplified by GPT-3 (175 billion parameters) (Brown et al., 2020) and more recently GPT-4 (Achiam et al., 2023), which has further expanded the boundaries of language understanding and generation. Similarly, vision-language models like CLIP (Radford et al., 2021) and Flamingo (Alayrac et al., 2022) have showcased the transformative power of scaling multimodal architectures. These advancements highlight how Scaling Up has enabled AI systems to achieve remarkable generalization and versatility across diverse tasks.
@@ 3.1 Reducing the Size of Large Models
Rather than modifying existing models, knowledge distillation (KD) facilitates the transfer of knowledge from large and complex teachers to small and efficient students (Hinton, 2015). The students are trained to replicate the behavior of the teachers. Yu et al. (2024) propose to distill System 2 reasoning processes—such as Chain-of-Thought and System 2 Attention—into a single-step System 1 model, eliminating intermediate reasoning while retaining or improving task performance. Program-aided Distillation (PaD) (Zhu et al., 2024) introduces a new KD paradigm that uses reasoning programs to verify and refine synthetic CoT data, enhancing distilled reasoning quality. PaD automates error-checking, incorporates iterative self-refinement to address faulty reasoning chains, and employs step-wise beam search to validate reasoning steps progressively.
@@ 3.3 Small Models for Large Impacts
More Parameters but Less Activation. A crucial trend in optimizing smaller models for efficiency is the adoption of Mixture-of-Experts (MoE) architectures, where a subset of model parameters is activated per token, reducing computation while maintaining a large overall parameter pool. Several recent models exemplify this technique: Mixtral (Jiang et al., 2024) consists of 8 expert models with a total of 56B parameters, while each token is processed by only 2 experts. Phi-3.5-MoE (Abdin et al., 2024) comprises 16 experts totaling 60.8B parameters but activates only 6.6B (10.9%). DeepSeek-V2 (Liu et al., 2024a) is a 236B-parameter model with 21B parameters activated per token (8.9%). DeepSeek-V3 (Liu et al., 2024b) scales further to 671B total parameters while activating only 37B per token (5.5%). These MoE-based models show that large parameter spaces can be efficiently leveraged without incurring excessive computational costs, allowing smaller models to achieve performance comparable to their larger counterparts with only a fraction of the active compute requirements.
@@ 4 Scaling Out: Advancing AI Ecosystems
We propose Scaling Out as the next step in AI evolution. Scaling Out expands AI’s reach by deploying interfaces—mechanisms that enable AI to interact with users, devices, and other systems. These interfaces, powered by specialized sub-models derived from foundation models, form an expandable AI ecosystem. Unlike Scaling Up’s focus on size or Scaling Down’s focus on efficiency, Scaling Out emphasizes accessibility and adaptability. For example, in a smart city, AI interfaces for traffic, energy, and safety could collaborate to create a seamless urban experience, showcasing Scaling Out’s transformative potential.
@@ 4.3 Future Trends
Edge Computing. Edge computing processes data locally on devices, such as smartphones, IoT sensors, or edge servers, minimizing the need to send information to centralized data centers. Federated learning complements this by allowing distributed devices to collaboratively train machine learning models without sharing raw data, enhancing privacy and efficiency. Together, these technologies reduce latency, improve real-time decision-making, and ensure scalability by distributing computation across a network of edge nodes. For Scaling Out, this decentralized architecture allows billions of lightweight, specialized AI agents to operate independently while sharing collective insights, as seen in applications like personalized healthcare monitoring or real-time traffic management. This synergy fosters ecosystems where agents adapt locally while contributing to a globally optimized intelligence network.
@@ 7 Alternative Views
While this position paper argues that Scaling Up encounters significant bottlenecks and that future trends will shift towards Scaling Down and Scaling Out, an alternative view is that Scaling Up remains a viable trajectory despite the challenges. Supporters argue that addressing these challenges is necessary and feasible through interdisciplinary innovation. Key challenges must be overcome include data quality, computational demands, and energy consumption.
Firstly, although data quantity has grown, their quality has not kept pace. Synthetic data offers a controlled alternative but may introduce biases and lack real-world applicability. As for computational demands, computational power is a bottleneck, as traditional hardware faces physical and economic limits. Alternatives like quantum, optical, and neuromorphic computing might help. Concern about energy consumption emphasizes the need for sustainable AI. Low-power chips and renewable energy integration could reduce the environmental impact of large-scale computing.
@@ 8 Conclusion
In this paper, we propose a framework for AI scaling, which is from Scaling Up to Scaling Down, then Scaling Out. Scaling Up lays the groundwork with foundation models that generalize across tasks. Scaling Down ensures efficiency and accessibility, optimizing AI for diverse environments. Finally, Scaling Out provides multiple AI interfaces, which enables collaborative intelligence and interaction with users to tackle real-world challenges.. Together, these advancements offer a vision of AI that amplifies human creativity, bridges societal divides, and empowers humanity to confront its most ambitious goals. The future of AI is not solely about technological breakthroughs, it is about building systems that are equitable, sustainable, and deeply integrated into the fabric of human progress.

### v2

# Position: AI Scaling: From Up to Down and Out

@@ 1 Introduction
The field of artificial intelligence (AI) has witnessed extraordinary advancements over the past decade, largely driven by the relentless pursuit of Scaling Up. Early breakthroughs were characterized by models with millions of parameters, such as AlexNet (Krizhevsky et al., 2012), word2vec (Church, 2017) and BERT (Devlin et al., 2019), which paved the way for deep learning’s success. This progression quickly escalated to models with billions of parameters, exemplified by GPT-3 (175 billion parameters) (Brown et al., 2020) and more recently GPT-4 (Achiam et al., 2023), which has further expanded the boundaries of language understanding and generation. Similarly, vision-language models like CLIP (Radford et al., 2021) and Flamingo (Alayrac et al., 2022) have showcased the transformative power of scaling multimodal architectures. These advancements highlight how Scaling Up has enabled AI systems to achieve remarkable generalization and versatility across diverse tasks.
@@
Notably, while our primary discussions in this paper center around LLMs, the proposed framework of Scaling Down and Scaling Out is not limited to them. The techniques and principles involved such as model compression, pruning and decentralized coordination are widely applicable across various AI systems, including vision models, reinforcement learning and multimodal architectures. By using LLMs as illustrative cases, we aim to ground the framework in a currently prominent domain, while maintaining its broader relevance.
@@ 3.1 Reducing the Size of Large Models
Rather than modifying existing models, knowledge distillation (KD) facilitates the transfer of knowledge from large and complex teachers to small and efficient students (Hinton, 2015). The students are trained to replicate the behavior of the teachers. Yu et al. (2024) propose to distill System 2 reasoning processes, such as Chain-of-Thought and System 2 Attention, into a single-step System 1 model, eliminating intermediate reasoning while retaining or improving task performance. Program-aided Distillation (PaD) (Zhu et al., 2024) introduces a new KD paradigm that uses reasoning programs to verify and refine synthetic CoT data, enhancing distilled reasoning quality. PaD automates error-checking, incorporates iterative self-refinement to address faulty reasoning chains, and employs step-wise beam search to validate reasoning steps progressively.
3.2 Reducing the Scale of Training
Recent efforts in dataset pruning have aimed to reduce training costs while preserving model performance, particularly on large-scale datasets. Traditional snapshot-based pruning methods estimate sample importance using static predictions from early or late training epochs but fail to generalize across architectures. To address these issues, He et al. (2023) propose Dynamic Uncertainty, which captures prediction variance over time using sliding windows, enabling better distinction between hard samples. Building on the need to incorporate temporal dynamics, the Temporal Dual-Depth Scoring (TDDS) (Zhang et al., 2024) method introduces a two-tiered scoring system that tracks both the stability and contribution of samples throughout training.
In the multimodal setting, Sieve (Mahmoud et al., 2024) addresses noise in web-crawled image-text pairs by using image captioning models and sentence transformers. YOCO (He et al., 2023) bridges the gap between dataset condensation and pruning by introducing Logit-Based Prediction Error and Balanced Construction that enable flexible resizing of condensed datasets without repeating the condensation process, offering a practical solution for constrained training environments. Together, these works reflect a shift toward dynamic, scalable, and generalizable data selection methods across both unimodal and multimodal domains.
@@ 3.4 Small Models for Large Impacts
More Parameters but Less Activation. A crucial trend in optimizing smaller models for efficiency is the adoption of MoE, where a subset of model parameters is activated per token, reducing computation while maintaining a large overall parameter pool. Several recent models exemplify this technique: Mixtral (Jiang et al., 2024) consists of 8 expert models with a total of 56B parameters, while each token is processed by only 2 experts. Phi-3.5-MoE (Abdin et al., 2024) comprises 16 experts totaling 60.8B parameters but activates only 6.6B (10.9%). DeepSeek-V2 (Liu et al., 2024a) is a 236B-parameter model with 21B parameters activated per token (8.9%). DeepSeek-V3 (Liu et al., 2024b) scales further to 671B total parameters while activating only 37B per token (5.5%).
@@ 4 Scaling Out: Advancing AI Ecosystems
We propose Scaling Out as the next step in AI evolution, which is a paradigm that leverages efficient, task-specific AI models derived from large-scale foundation models, distributed across networks and interfacing through modular, interactive systems. Scaling Out can expand AI’s reach by deploying interfaces, which enable AI to interact with users, devices, and other systems. These interfaces, powered by specialized sub-models derived from foundation models, form an expandable AI ecosystem. Unlike Scaling Up’s focus on size or Scaling Down’s focus on efficiency, Scaling Out emphasizes accessibility and adaptability. For example, in a smart city, AI interfaces for traffic, energy, and safety could collaborate to create a seamless urban experience, showcasing Scaling Out’s transformative potential.
@@ 4.1 Scaling Out builds an AI Ecosystem
To make this concept tangible, we provide some analogies and real-world illustrations. (i) In a functioning society, individuals often specialize in different roles like teachers, policemen, and engineers. Each adapts to their context while contributing to maintaining social order and driving progress. For example, a teacher adjusts their methods to meet the needs of students, while contributing to the broader educational infrastructure. This mirrors how Scaling Out in AI enables diverse, task-specific models to operate within their domains, while remaining connected through structured interfaces. (ii) In healthcare systems, when diagnosing and treating a patient, hospitals do not rely on a single generalist. Instead, responsibilities are distributed across specialized professionals such as GPs handling basic diagnostics, radiologists interpreting medical images, surgeons performing targeted procedures and pharmacists managing prescriptions. Each operates based on their expertise and the specific contextual information they receive. These experts interact through structured protocols (e.g., referrals), and their collaboration results in an adaptive, efficient, and context-aware response tailored to individual needs. This reflects how AI systems under Scaling Out can coordinate modular intelligence across specialized sub-models.
@@ 4.2 Technical Foundations
Protocol and System-Level Coordination. As Scaling Out increasingly depends on collaboration among specialized models and agents, standard communication protocols play a key role. Google’s recent Agent-to-Agent (A2A)222https://github.com/google/A2A initiative introduces an open protocol that enables agents built on different frameworks to interoperate, negotiate interaction formats, and securely collaborate. Similarly, the emerging Model Context Protocol (MCP)333https://modelcontextprotocol.io/introduction provides a standardized interface for model-to-model interactions. MCP facilitates consistent context sharing, input/output formatting, and execution state management across diverse models.
@@ 4.3 Future Trends
Edge Computing. Edge computing processes data locally on devices, such as smartphones, IoT sensors, or edge servers, minimizing the need to send information to centralized data centers. Federated learning complements this by allowing distributed devices to collaboratively train models without sharing raw data. Together, these technologies reduce latency, improve real-time decision-making, and ensure scalability by distributing computation across a network of edge nodes. For Scaling Out, this decentralized architecture allows billions of lightweight, specialized AI agents to operate independently while sharing collective insights, as seen in applications like personalized healthcare monitoring or real-time traffic management. This synergy fosters ecosystems where agents adapt locally while contributing to a globally optimized intelligence network.
@@ 7 Alternative Views
While this position paper argues that Scaling Up encounters significant bottlenecks and that future trends will shift towards Scaling Down and Scaling Out, an alternative view is that Scaling Up remains a viable trajectory despite challenges. Supporters argue that addressing challenges of Scaling Up is necessary and feasible through interdisciplinary innovation. Key challenges must be overcome include data quality, computational demands, and energy consumption.
Firstly, although data quantity has grown, their quality has not kept pace. Synthetic data offers a controlled alternative but may introduce biases and lack real-world applicability. As for computational demands, computational power is a bottleneck, as traditional hardware faces physical and economic limits. Alternatives like quantum, optical, and neuromorphic computing might help. Concern about energy consumption emphasizes the need for sustainable AI. Low-power chips and integration of renewable energy could reduce the environmental impact of large-scale computing.
In addition, the advantage of increasing the scale of AI models also lies in that it can achieve significant advancements in their reasoning and generalization capabilities. Furthermore, Scaling Up has enabled transformative progress in multi-modal generative AI systems, enhancing the quality and fidelity of generated images and videos which smaller models often struggle to match.
@@ 8 Conclusion
In this paper, we propose a framework for AI scaling, which is from Scaling Up to Scaling Down, then Scaling Out. Scaling Up lays the groundwork with foundation models that generalize across tasks. Scaling Down ensures efficiency and accessibility, optimizing AI for diverse environments. Finally, Scaling Out provides multiple AI interfaces, which enables collaborative intelligence and interaction with users to tackle real-world challenges. Together, these advancements offer a vision of AI that amplifies human creativity, bridges societal divides, and empowers humanity to confront its most ambitious goals. The future of AI is not solely about technological breakthroughs, it is about building systems that are equitable, sustainable, and deeply integrated into the fabric of human progress.
Evaluating Scaling Down and Out requires going beyond accuracy to include efficiency metrics like FLOPs, latency, and energy use. For Scaling Down, metrics such as cost-per-inference, performance-per-watt, and capability-per-dollar better reflect practical value. For Scaling Out, evaluation should consider ecosystem-level indicators such as the scalability of specialized models, robustness of distributed deployments, and diversity of fine-tuned models on open platforms such as Hugging Face or OpenAI’s GPT store.
Acknowledgements
This work was supported in part by the Australian Research Council under Projects DP240101848 and FT230100549.